# Phase transition on the fluctuation of the structure of random walk ranges

### v1

@@ 3.2. Proof of CLT
$\|\overline{X}_{2n}\|_{2+\epsilon}\leq\|\overline{X}_{n}+\overline{X}[n,2n]\|_{2+\epsilon}+\|E_{n}\|_{2+\epsilon},$
@@ 4.1.1. Graph distance and cut points
We define the error term
$\mathcal{E}_{k}:=\sum_{l=1}^{2^{k}}-E_{n}^{(k,l)}.$
@@
$\mathrm{Var}(\mathcal{E}_{k})=\sum_{l=1}^{2^{k}}\mathrm{Var}(E_{n}^{(k,l)})\lesssim\sum_{l=1}^{2^{k}}\frac{n}{2^{k}}=n.$
Thus we have, by expanding the variance and applying the Cauchy-Schwarz inequality, that
@@
In addition, since each $X[k,k+1]$ is deterministically 1, we have that $\mathrm{Var}(X[k,k+1])=0$. Then, we have
@@
$\mathbb{E}[E_{n}^{(k_{1},l_{1})}E_{n}^{(k_{2},l_{2})}]\lesssim(\log n)^{2}.$
Let $\tilde{S}$ be the reversal of the first part, that is $\tilde{S}_{i}^{(k,l)}=S_{l\frac{n}{2^{k}}-i}-S_{l\frac{n}{2^{k}}}$ . $\hat{S}$ will be the second part, that is $\hat{S}_{i}^{(k,l)}=S_{l\frac{n}{2^{k}}+i}-S_{l\frac{n}{2^{k}}}$. Indeed, if $E_{n}^{(k,l)}\geq a$, $\Omega(a,k,l)$ holds, where
@@
$\leq$ $\mathbb{P}\bigg(\{\hat{S}^{(k_{1},l_{1})}[\frac{a}{2},\frac{n}{2^{k_{1}}}]\cap\tilde{S}^{(k_{1},l_{1})}[0,\frac{n}{2^{k_{1}}}]\neq\emptyset\}\cap\{\hat{S}^{(k_{2},l_{2})}[\frac{b}{2},\frac{n}{2^{k_{2}}}]\cap\tilde{S}^{(k_{2},l_{2})}[0,\frac{n}{2^{k_{2}}}]\neq\emptyset\}\bigg)$
@@
Therefore,
$\mathrm{Var}\left(\sum_{k=0}^{a_{n}}\mathcal{E}_{k}\right)\leq\sum_{k=0}^{\log_{2}n}\mathrm{Var}(\mathcal{E}_{k})+2\sum_{k_{1}=0}^{\log_{2}(a_{n})}\sum_{l_{1}=1}^{2^{k_{1}}}\sum_{k_{2}=k_{1}+1}^{\log_{2}(a_{n})}\sum_{l_{2}=0}^{2^{k_{2}-k_{1}}}\mathbb{E}[E_{n}^{(k_{1},l_{1})}E_{n}^{(k_{2},l_{2})}]\lesssim n(\log n)$
@@
Fix $\delta>0$ very close to 1. We will see where this condition plays a role close to the end of the proof.
By our binary decomposition, we find that,
@@
$\mathrm{Var}(\mathcal{E})=\sum_{k=0}^{\log_{2}(n^{\delta})}\sum_{l=1}^{2^{k}}\mathrm{Var}(-E_{n}^{(k,l)})+\sum_{k_{1}=0}^{\log_{2}(n^{\delta})}\sum_{l_{1}=1}^{2^{k_{1}}}\sum_{k_{2}=0}^{\log_{2}(n^{\delta})}\sum_{l_{2}=1}^{2^{k_{2}}}\text{Cov}(-E_{n}^{(k_{1},l_{1})},-E_{n}^{(k_{2},l_{2})}).$
@@
Furthermore, we remark that if the intervals $n/2^{k_{1}+1}(2l_{1}-2),n/2^{k_{1}+1}(2l_{1})]$ (this is the interval corresponding to the cross term $E^{(k_{1},l_{1})}$) and $n/2^{k_{2}+1}(2l_{2}-2),n/2^{k_{2}+1}(2l_{2})]$(this is the interval corresponding to $E^{(k_{2},l_{2})}$) are completely disjoint from each other, then $E^{(k_{1},l_{1})}$ and $E^{(k_{2},l_{2})}$ are independent of each other and the covariance is 0. Using these considerations, we see that a lower bound for the covariances is the following:
$\sum_{k_{1}=0}^{\log_{2}(n^{\delta})}\sum_{l_{1}=1}^{2^{k_{1}}}\sum_{k_{2}=k_{1}+1}^{\log_{2}(n^{\delta})}\sum_{l_{2}=0}^{2^{k_{2}-k_{1}}}-(\log n)^{2}\geq-\sum_{k_{1}=0}^{\log_{2}(n^{\delta})}2^{k_{1}}[n^{\delta}/2^{k_{1}}](\log n)^{2}$
@@
by setting $\delta$ sufficiently close to $1$, this can be made very small.
@@
Then for $\delta$ close to but not equal to $1$, we have $\mathrm{Var}(\mathcal{E})\geq cn\log n$ and $\mathrm{Var}(MT)\leq C(1-\delta)n\log n$. Thus, $\mathrm{Var}(X_{n})\gtrsim n\log n$, as desired. ∎
4.1.2. Effective resistance
We start with the following lower bound.
@@ 4.1.2. Effective resistance
Fix $\delta>0$. For $\delta>0$, we will choose a sufficiently small $\kappa>0$ and derive a contradiction later. Recall our binary decomposition in (4.2). As long as $n^{1-\delta}\geq N_{\kappa}$, we have that
(4.3) $\mathrm{Var}(MT)\leq n^{\delta}[\kappa n^{1-\delta}\log n^{1-\delta}]=\kappa(1-\delta)n\log n$
and
We have that, for sufficiently large $n$,
Notice that
$\text{Cov}(-E_{n}^{(k_{1},l_{1})},-E_{n}^{(k_{2},l_{2})})=\mathbb{E}[E_{n}^{(k_{1},l_{1})}E_{n}^{(k_{2},l_{2})}]-\mathbb{E}[E_{n}^{(k_{1},l_{1})}]\mathbb{E}[E_{n}^{(k_{2},l_{2})}]$
$\geq-\mathbb{E}[E_{n}^{(k_{1},l_{1})}]\mathbb{E}[E_{n}^{(k_{2},l_{2})}]\gtrsim-(\log n)^{2}.$
Here, we only used that $E_{n}\geq 0$ to get the second inequality.
Putting these estimates back into equation (4.4), we observe that
(4.5) $\mathrm{Var}(\mathcal{E})\geq c\delta n\log n-O((\log n)^{2}(n^{\delta}\log(n^{\delta}))^{2})$
This last term will be greater than $\frac{c\delta}{2}n\log n$ for $n$ sufficiently large.
Finally, we can apply the Cauchy-Schwarz inequality to say that,
@@
Notice that as a function of the value of $\mathrm{Var}(\mathcal{E})$, the right-hand side is increasing when $\mathrm{Var}(\mathcal{E})\geq\mathrm{Var}(MT).$ We have already obtained a lower bound on $\mathrm{Var}(\mathcal{E})$ from equation (4.5) that is larger than the upper bound we have on $\mathrm{Var}(MT)$ from equation (4.3).
Indeed, notice that if $\mathrm{Var}(MT)<\frac{c\delta}{8}n\log n$, then we would have that the right hand side of (4.6) will be greater than,
@@
Then, we have that by setting $\kappa$ sufficiently small that and applying the previous inequality that $\mathrm{Var}(X_{n})\geq\frac{c\delta}{8}n\log n$. This is a contradiction to the assumption that we cannot produce a lower bound to the $\limsup$ of $\frac{\mathrm{Var}(X_{n})}{n\log n}$. ∎
Now, we produce an upper bound to the variance of the effective resistance. Since the proof is the same as that of (4.1), we omit it.
@@ 4.2.1. Graph distance and cut points
The variables $Z_{n_{\delta},i}$ are all independent of each other and one
@@
(4.7) $\mathbb{P}(\mathcal{M}\geq\sqrt{n(\log n)^{1-\delta/2}})\to 0$
@@
Recall that $\mathcal{G}$ from Part 1 was a Gaussian random variable and the order of the variance of $\mathcal{G}$ was $n\log n$, which is much larger than the error terms of Parts 2 and 3. Then, combining Part 1, 2, 3, we can obtain the desired result using the Lindeberg–Feller central limit theorem.
4.2.2. Effective resistance
Next, we will give the proof of Theorem 1.4 for $d=6$. We present only the necessary adjustments from the proof of Theorem 1.2. This involves changing Parts 1 and 4, while keeping Parts 2 and 3. We also remark that the results of this section only depend on the tail behavior of the cross terms $E_{n}$ behaving like $\mathbb{P}(E_{n}\geq l)\asymp\frac{1}{l}$. In particular, the results are general enough to not require any estimate on the correlation between the cross terms. Thus, the methods of this section are of general interest and can potentially be applied to other heavy tailed statistics of the random walk range.
@@ 4.2.2. Effective resistance
for all $k$ large enough. This will imply that $\lim_{k\to\infty}f(m_{k})=\infty$. This contradicts the upper bound on $f$ from Lemma 4.1.
Thus, we see that this proof establishes that there is a subsequence where $\frac{X_{n}-\mathbb{E}[X_{n}]}{\sqrt{\mathrm{Var}(X_{n})}},$ does not converge to 0. Moreover, along this subsequence $X_{n}-\mathbb{E}[X_{n}]$ will be distributed (after appropriate normalization) as a normal random variable whose variance is given by $n(\log(n/n_{\delta}))^{f(n/n_{\delta})}$. ∎
@@ 5. The fluctuations converge to 0 in dimension $d=5$
In this section, we will give the proof of Propositions 1.1 and 1.3 and Theorems 1.2 and 1.4 in $d=5$. Since we do not show the correlation estimates (5.1) for the effective resistance, we present the weaker estimates of its variance. Since the proof is similar, we will omit the proof of Proposition 1.3 in $d=5$.
@@
Hence, we obtain $\mathrm{Var}(\mathcal{E})\gtrsim n^{3/2}$.
In addition, (2.2) yields $\sum_{k=1}^{(1-\alpha)\log_{2}n}\sum_{l=1}^{2^{k}}\mathbb{E}[(E_{n}^{(k,l)})^{2}]\leq cn^{3/2}$. Now by the same argument as that of Lemma 4.1, we have
@@
Proof of Theorems 1.2 and 1.4 in $d=5$.
For small $\delta>0$, we use Markov’s inequality in the computation of the variance, to derive,
@@
Now, we use Markov’s inequality in the computation of the absolute value combined with the triangle inequality, to get,
@@ 6. The fluctuations converge to 0 in dimension $d=4$
In this section, we estimate the variance and the Limit in distribution of $\frac{X_{n}-\mathbb{E}[X_{n}]}{\sqrt{\mathrm{Var}(X_{n})}}$ in $d=4$.
In what follows, we will simplify notation by reducing to a discussion of only the cross term $E_{n}$ since we will only apply a dyadic decomposition and the only other cross term that could appear, $X_{2n+1}-X_{n}-X[n,2n+1]$, will have a similar analysis.
In subsection 6.1, we will give the lower bounds for the cross term, which holds for the graph distance, cut points and the effective resistance. In subsections 6.2-6.4, we will show the upper bounds for the cross term and main results in $d=4$. In these subsections, we are considering only the graph distance, cut points since Lemma 6.8 does not work for the effective resistance. We will give our main result in $d=4$ for the effective resistance in subsection 6.5.

### v2

@@ 3.2. Proof of CLT
$\|\overline{X}_{2n}\|_{2+\epsilon}\leq\|\overline{X}_{n}+\overline{X}[n,2n]\|_{2+\epsilon}+\|E_{n}\|_{2+\epsilon}.$
@@
Remark 3.6.
We remark here that all the proofs in this section will hold without change as long as we assume the tail statistic $\mathbb{P}(X_{n}+X_{[n,2n]}-X_{2n}\geq l)\leq Cl^{-d/2+2}$, for $d\geq 7$. This establishes Proposition 1.4 and Theorem 1.5 for $d\geq 7$.
@@ 4.1.1. Graph distance and cut points
We define
$\mathcal{E}_{k}:=\sum_{l=1}^{2^{k}}-E_{n}^{(k,l)},\quad\mathcal{E}:=\sum_{k=0}^{\log_{2}n}\mathcal{E}_{k},\quad MT:=\sum_{k=0}^{n}X[k,k+1].$
@@
(4.1) $\mathrm{Var}(\mathcal{E}_{k})=\sum_{l=1}^{2^{k}}\mathrm{Var}(E_{n}^{(k,l)})\lesssim\sum_{l=1}^{2^{k}}\frac{n}{2^{k}}=n.$
By expanding the variance and applying the Cauchy-Schwarz inequality, we obtain
@@
In addition, since each $X[k,k+1]$ is deterministically 1, we have that $\mathrm{Var}(X[k,k+1])=0$. Then,
@@
Furthermore, we remark that if the intervals $n/2^{k_{1}+1}(2l_{1}-2),n/2^{k_{1}+1}(2l_{1})]$ (, which is the interval corresponding to the cross term $E^{(k_{1},l_{1})}$) and $n/2^{k_{2}+1}(2l_{2}-2),n/2^{k_{2}+1}(2l_{2})]$(this is the interval corresponding to $E^{(k_{2},l_{2})}$) are completely disjoint, then $E^{(k_{1},l_{1})}$ and $E^{(k_{2},l_{2})}$ are independent and the covariance is 0. Thus, the only non-zero covariances $\text{Cov}(E^{(k_{1},l_{1})},E^{(k_{2},l_{2})})$ are those with $k_{2}\geq k_{1}$ and $2^{k_{2}-k_{1}}l_{1}\leq l_{2}\leq 2^{k_{2}-k_{1}}(l_{1}+1)$.
@@
$\mathbb{E}[E_{n}^{(k_{1},l_{1})}E_{n}^{(k_{2},l_{2})}]\lesssim(\log n)^{2},$
for most pairs of $(k_{1},l_{1})$ and $(k_{2},l_{2})$ with non-trivial intersection. Let $\tilde{S}$ be the reverse of the first part, that is $\tilde{S}_{i}^{(k,l)}=S_{l\frac{n}{2^{k}}-i}-S_{l\frac{n}{2^{k}}}$ . $\hat{S}$ will be the second part, that is $\hat{S}_{i}^{(k,l)}=S_{l\frac{n}{2^{k}}+i}-S_{l\frac{n}{2^{k}}}$. Indeed, if $E_{n}^{(k,l)}\geq a$, $\Omega(a,k,l)$ holds, where
@@
$\leq$ $\mathbb{P}\bigg(\{\hat{S}^{(k_{1},l_{1})}\left[\frac{a}{2},\frac{n}{2^{k_{1}}}\right]\cap\tilde{S}^{(k_{1},l_{1})}\left[0,\frac{n}{2^{k_{1}}}\right]\neq\emptyset\}\cap\{\hat{S}^{(k_{2},l_{2})}\left[\frac{b}{2},\frac{n}{2^{k_{2}}}\right]\cap\tilde{S}^{(k_{2},l_{2})}\left[0,\frac{n}{2^{k_{2}}}\right]\neq\emptyset\}\bigg)$
@@
Therefore, with the aid of (4.1), we obtain,
(4.3) $\mathrm{Var}\left(\sum_{k=0}^{a_{n}}\mathcal{E}_{k}\right)$
$\leq$ $\sum_{k=0}^{\log_{2}n}\mathrm{Var}(\mathcal{E}_{k})+2\sum_{k_{1}=0}^{\log_{2}(a_{n})}\sum_{l_{1}=1}^{2^{k_{1}}}\sum_{k_{2}=k_{1}+1}^{\log_{2}(a_{n})}\sum_{l_{2}=2^{k_{2}-k_{1}}l_{1}}^{2^{k_{2}-k_{1}}(l_{1}+1)}\text{Cov}(E_{n}^{(k_{1},l_{1})},E_{n}^{(k_{2},l_{2})})$
$\leq$ $\sum_{k=0}^{\log_{2}n}\mathrm{Var}(\mathcal{E}_{k})+2\sum_{k_{1}=0}^{\log_{2}(a_{n})}\sum_{l_{1}=1}^{2^{k_{1}}}\sum_{k_{2}=k_{1}+1}^{\log_{2}(a_{n})}\sum_{l_{2}=2^{k_{2}-k_{1}}l_{1}}^{2^{k_{2}-k_{1}}(l_{1}+1)}\mathbb{E}[E_{n}^{(k_{1},l_{1})}E_{n}^{(k_{2},l_{2})}]\lesssim n(\log n)$
@@
Fix $\delta>0$ very close to 1. We will see where this condition comes into play near the end of the proof. By our binary decomposition, we find that,
@@
$\sum_{k_{1}=0}^{\log_{2}(n^{\delta})}\sum_{l_{1}=1}^{2^{k_{1}}}\sum_{k_{2}=k_{1}+1}^{\log_{2}(n^{\delta})}\sum_{l_{2}=2^{k_{2}-k_{1}}l_{1}}^{2^{k_{2}-k_{1}}(l_{1}+1)}-(\log n)^{2}\geq-\sum_{k_{1}=0}^{\log_{2}(n^{\delta})}2^{k_{1}}[n^{\delta}/2^{k_{1}}](\log n)^{2}$
@@
By combining these estimates, we have,
(4.6) $\mathrm{Var}(\mathcal{E})\geq c\delta n\log n.$
@@
setting $\delta$ sufficiently close to $1$, this can be made very small.
@@
Then for $\delta$ close to but not equal to $1$, we have $\mathrm{Var}(\mathcal{E})\geq c\delta n\log n$ and $\mathrm{Var}(MT)\leq C(1-\delta)n\log n$. Thus, $\mathrm{Var}(X_{n})\gtrsim n\log n$, as desired. ∎
4.1.2. General Analysis with heavy tailed Cross Terms
In this section, we will just assume that we are considering a collection of variables $X_{[a,b]}$ that are measurable with respect to the walk $\{S_{a},\ldots,S_{b}\}$ and that the cross term $E_{n}:=X_{n}+X_{[0,n]}-X_{2n}$ satisfies the tail bounds $\mathbb{P}(E_{n}\geq l)\asymp\frac{1}{l}$. We start with the following lower bound.
@@ 4.1.2. General Analysis with heavy tailed Cross Terms
Fix $\delta>0$. For $\delta>0$, we will choose a sufficiently small $\kappa>0$ and derive a contradiction later. Recall our binary decomposition in (4.4).
As long as $n^{1-\delta}\geq N_{\kappa}$, we have that
$\mathrm{Var}(MT)\leq n^{\delta}[\kappa n^{1-\delta}\log n^{1-\delta}]=\kappa(1-\delta)n\log n.$
In addition, equation (4.6), will give us $\text{Var}(\mathcal{E})\geq c\delta n\log n.$ Finally, we can apply the Cauchy-Schwarz inequality to say that,
@@
Notice that, as a function of the value of $\mathrm{Var}(\mathcal{E})$, the right-hand side increases when $\mathrm{Var}(\mathcal{E})\geq\mathrm{Var}(MT)$. Indeed, notice that if $\mathrm{Var}(MT)<\frac{c\delta}{8}n\log n$, then the right hand side of (4.7) will be greater than,
@@
Then, by setting $\kappa$ sufficiently small and applying the previous inequality, we have that $\mathrm{Var}(X_{n})\geq\frac{c\delta}{8}n\log n$. This is a contradiction to the assumption that we cannot produce a lower bound to the $\limsup$ of $\frac{\mathrm{Var}(X_{n})}{n\log n}$. ∎
Now, we produce an upper bound to the variance of general variables with the same heavy tail on the cross terms. Since the proof is the same as that of (4.2), we omit it.
@@ 4.2.1. Graph distance and cut points
The variables $Z_{n_{\delta},i}$ are all independent of each other and one obtains
@@
(4.8) $\mathbb{P}\bigg(\mathcal{M}\geq\sqrt{n(\log n)^{1-\delta/2}}\bigg)\to 0$
@@
Recall that $\mathcal{G}$ from Part 1 was a Gaussian random variable, and the order of the variance of $\mathcal{G}$ was $n\log n$, which is much larger than the error terms of Parts 2 and 3. Then, combining Parts 1, 2, and 3, we can obtain the desired result using the Lindeberg–Feller central limit theorem.
4.2.2. General Analysis with heavy tailed Cross Terms
In this section, we will demonstrate a version of the phase transition demonstrated for the graph distance of the random walk range in six dimensions for variables whose cross terms have the same heavy tailed asymptotics: $\mathbb{P}(E_{n}\geq l)\asymp l^{-1}$. In particular, we do not make any assumptions about the covariance structure for the cross terms. We emphasize that the results of this section show that heavy tailed behavior is the key source of the phase transition. We note that our results will apply to the effective resistance; we briefly give the definitions here.
This corresponds to the energy dissipation of an electrical network on the graph $G$. We also introduce the space
$H^{2}:=\{\,f\in\mathbb{R}^{V}:\mathcal{E}(f,f)<\infty\,\},$
This quantity represents the voltage difference required to send the current from $A$ to $B$, and therefore corresponds to the usual notion of effective resistance in an electrical network. For vertices $x,y\in V$, we write $R_{G}(x,y)$ to denote $R_{G}(\{x\},\{y\})$.
Now, we will give the proof of Theorem 1.5 for $d=6$. We present only the necessary adjustments from the proof of Theorem 1.2. This involves changing Parts 1 and 4, while keeping Parts 2 and 3. As before, we mention that the particular value of this result is that it demonstrates that the right order of heavy tails is what ultimately leads to a phase transition in dimension 6, rather than detailed estimates of the covariance structure.
@@ 4.2.2. General Analysis with heavy tailed Cross Terms
for all $k$ large enough. This will imply that $\lim_{k\to\infty}f(m_{k})=\infty$. This contradicts the upper bound of $f$ from Lemma 4.1.
Thus, we see that this proof establishes that there is a subsequence where $\frac{X_{n}-\mathbb{E}[X_{n}]}{\sqrt{\mathrm{Var}(X_{n})}},$ does not converge to 0. Moreover, along this subsequence, $X_{n}-\mathbb{E}[X_{n}]$ will be distributed (after appropriate normalization) as a normal random variable whose variance is given by $n(\log(n/n_{\delta}))^{f(n/n_{\delta})}$. ∎
@@ 5. The fluctuations converge to 0 in dimension $d=5$
In this section, we will give the proof of Propositions 1.1 and 1.4 and Theorems 1.2 and 1.5 in $d=5$. Since the proof is similar, we will omit the proof of Proposition 1.4 in $d=5$.
@@
Hence, we obtain $\mathrm{Var}(\mathcal{E})\gtrsim n^{3/2}$. In addition, (2.2) yields $\sum_{k=1}^{(1-\alpha)\log_{2}n}\sum_{l=1}^{2^{k}}\mathbb{E}[(E_{n}^{(k,l)})^{2}]\leq cn^{3/2}$. Now by the same argument as that of Lemma 4.1, we have
@@
Proof of Theorems 1.2 and 1.5 in $d=5$.
For small $\delta>0$, we use Markov’s inequality in the computation of the variance to derive,
@@
Now, we use Markov’s inequality, along with the triangle inequality, to get,
@@ 6. The fluctuations converge to 0 in dimension $d=4$
In this section, we estimate the variance and the limit in distribution of $\frac{X_{n}-\mathbb{E}[X_{n}]}{\sqrt{\mathrm{Var}(X_{n})}}$ in $d=4$. In subsection 6.1, we will give the lower bounds for the cross term, which holds for the graph distance and the cut points. In subsections 6.2-6.4, we will show the upper bounds for the cross term and main results in $d=4$. In subsection 6.5, we will present our results for general random variables whose cross terms have the same heavy tailed behavior as the graph distance in dimension 4.